# references.bib

### v1

@@ IV Main Results
In this section, under the proper assumptions, we will show that $A^{*}$ is a solution to the $l_{1}$-norm estimator after a finite time. Let $\hat{A}_{T}$ denote any estimate obtained from (11), which can be equivalently written as
@@ IV-A Conditions for the Exact Recovery
In this subsection, we provide necessary and sufficient conditions for the exact recovery. For the following theorem, let $x_{t}^{i}$ and $w_{t}^{i}$ denote the $i^{\text{th}}$ entries of $x_{t}$ and $w_{t}$, respectively. Define the random variables $z_{t}^{ij+}$ and $z_{t}^{ij-}$ as follows:
[…]
Theorem 1
$A^{*}$ is a solution to the $l_{1}$-norm estimator (11) at time $T$ if and only if
$\sum_{t=0}^{T-1}z_{t}^{ij+}\geq 0,$ (14a)
$\sum_{t=0}^{T-1}z_{t}^{ij-}\geq 0$ (14b)
for all $i,j\in\{1,\dots,d\}$.
Proof:
Since the non-differential minimization problem (12) is convex and unconstrained, the Karush-Kuhn-Tucker (KKT) conditions provide the necessary and sufficient conditions for optimality:
$0\in\sum_{t=0}^{T-1}\partial\|(A^{*}-\hat{A}_{T})x_{t}+w_{t}\|_{1}\cdot x_{t}^{T}.$ (15)
Thus, $A^{*}$ is a solution to (12) (or equivalently (11)) if and only if
$0\in\sum_{t=0}^{T-1}\partial\|w_{t}\|_{1}\cdot x_{t}^{T}.$ (16)
Note that both sides of (16) are $d\times d$ matrices. Using Definition 4, the equivalent form of (16) is
$-\sum_{t:w_{t}^{i}=0}|x_{t}^{j}|\leq-\sum_{t:w_{t}^{i}<0}x_{t}^{j}+\sum_{t:w_{t}^{i}>0}x_{t}^{j}\leq\sum_{t:w_{t}^{i}=0}|x_{t}^{j}|$ (17)
for all $i,j\in\{1,\dots,d\}$. One can separate (17) into two inequalities and attain $2d^{2}$ number of inequalities, which are again rearranged to (14a) and (14b). ∎
To attain the exact recovery, the inequalities in (14) must hold. Since the left-hand sides of (14a) and (14b) are random variables, it is necessary to study their expected values and ensure that they are sufficiently positive with a high probability. We require a mild assumption stating that no adversarial disturbance can deceive the system to the extent that each coordinate of the next state collapses to absolute zero, which ensures sufficient exploration of the state trajectory.
Assumption 4 (Non-degeneracy)
There exists $\lambda>0$ such that
$\mathbb{E}[|x_{0}^{i}|]\geq\lambda,\quad$ $i=1,\dots,d,$ (18a)
$\mathbb{E}[|x_{t+1}^{i}|~{}|~{}\mathcal{F}_{t}]\geq\lambda,\quad$ $i=1,\dots,d$ (18b)
for all $t$ and $\mathcal{F}_{t}$.
In the following two subsections, each will present a distinct scenario in which the $l_{1}$-norm estimator achieves the exact recovery despite the presence of adversarial noncentral disturbances.
@@ IV-B Exact Recovery under Sign-restricted Disturbances
In this subsection, we will demonstrate that the exact recovery is achieved under the condition that each disturbance has equal chance of being positive or negative. The structure of the disturbance $w_{t}$ consists of $\alpha_{t}$ and $\beta_{t}$, where $\alpha_{t}$ is a sub-Gaussian vector symmetric around zero and $\beta_{t}>0$ is an amplifying factor. The formal assumption is given below.
Assumption 5 (Sign-restricted disturbances)
Given the filtration $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$, a disturbance $w_{t}\in\mathbb{R}^{d}$ is formulated as
@@
  • $\beta_{t}$ is an amplifying random vector conditioned on $\mathcal{F}_{t}$, where for all $i=1,\dots,d$, there exists $B>0$ such that $0<\beta_{t}^{i}\leq B$.
@@
Remark 1 (Necessity of symmetry assumption)
To formulate adversarial noncentral disturbances, we have introduced $\alpha_{t}$ that is symmetric around zero. We adopted this assumption since the system cannot defend the attack if the adversary has a full information set $\mathcal{F}_{t}$ and the freedom of choosing any nonzero-mean attack. For example, consider a scalar system $x_{t+1}=a^{*}x_{t}+w_{t}$, where the true system satisfies $10<a^{*}<$ 1 and $10<x_{0}<$ 1. Suppose that $w_{t}$ conditioned on $\mathcal{F}_{t}$ is $-\text{sgn}(x_{t})$ for all $t=0,1,\dots,$ where $\text{sgn}(\cdot)$ is defined in (8). In such a case, one always attains
@@
However, any valid estimation method tries to minimize the distance between $x_{t+1}$ and $\hat{a}x_{t}$ to find an estimate $\hat{a}$. One can thus only arrive at a negative estimate $\hat{a}$ despite the true system $a^{*}>0$, since we have $\text{sgn}(x_{t})\neq\text{sgn}(x_{t+1})$ for all $t=0,1,\dots$. This occurs since the attacker can always deceive the system if the attack can be universally positive or negative depending on the filtration. The symmetry assumption (20) on $\alpha_{t}$ prevents such a phenomenon and enables system identification.
The disturbance structure (19) includes zero-mean Gaussian random variables since $\alpha_{t}$ can be zero-mean Gaussian distribution and $\beta_{t}$ can be a constant random vector. Our formulation generalizes this by converting a symmetric distribution into a noncentral distribution with a random vector $0<\beta_{t}<\infty$.
Now, we analyze the conditions for the exact recovery shown in Theorem 1. The following lemma provides the expectation information of (14).
Lemma 3
Under Assumptions 3, 4, and 5, we have
$\mathbb{E}\biggr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\biggr{]}\geq(1-p)\lambda T,$ (21a)
$\mathbb{E}\biggr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\biggr{]}\geq(1-p)\lambda T.$ (21b)
Proof:
Note that the symmetry assumption (20) implies that the adversarial disturbance $w_{t}$ also satisfies
@@
for all $\mathcal{F}_{t}$ due to $\beta_{t}>0$. By the definition of $z_{t}^{ij+}$ given in (13), we have
$\mathbb{E}$ $\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\Bigr{]}=\sum_{t=0}^{T-1}\mathbb{E}[z_{t}^{ij+}]$
$=\sum_{t=0}^{T-1}\mathbb{E}\bigr{[}|x_{t}^{j}|\cdot\mathbb{I}(w_{t}^{i}=0)+x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}<0)-x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}>0)\bigr{]}$
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}\mathbb{E}\bigr{[}|x_{t}^{j}|\cdot\mathbb{I}(w_{t}^{i}=0)$
$\hskip 56.9055pt+x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}<0)-x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}>0)~{}\bigr{|}~{}\mathcal{F}_{t}]\Bigr{]}$
@@
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}|x_{t}^{j}|\cdot\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})$
$\hskip 33.57425pt+x_{t}^{j}\cdot[\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})]\Bigr{]}$ (22)
$\geq\sum_{t=0}^{T-1}(1-p)\cdot\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}+0\geq(1-p)\lambda T,$
due to the tower rule and the fact that the filtration $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$ contains the information of $x_{t}$. The last inequality is due to Assumptions 3 and 4. Similarly, $\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\bigr{]}$ has the same lower bound. ∎
Now, it remains to bound the tail probability of the left-hand sides of (14a) and (14b), for which we will study the sub-Gaussian norms.
Lemma 4
Under Assumptions 1 and 2, we have
$\mathbb{E}\biggr{[}\exp\biggr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\biggr{)}\biggr{]}$
@@
$\hskip 85.35826pt\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)},$ (23a)
@@
$\mathbb{E}\biggr{[}\exp\biggr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\bigr{]}\Bigr{]}\biggr{)}\biggr{]}$

### v5

@@ IV Main Results
In this section, under proper assumptions, we will show that the $l_{1}$-norm estimator achieves the finite-time exact recovery defined in Definition 4. Let $\hat{A}_{T}$ denote any estimate obtained from (11), which can be equivalently written as
@@ IV-A Conditions for the Exact Recovery
In this subsection, we provide sufficient conditions for the exact recovery. For the following theorem, let $w_{t}^{i}$ denote the $i^{\text{th}}$ entry of $w_{t}$. Given $y\in\mathbb{R}^{d}$, define the random variables $z_{t}^{i}(y)$ as follows:
$z_{t}^{i}(y)=\begin{dcases*}|y^{T}x_{t}|,&if $w_{t}^{i}=0$,\\ y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i}),&otherwise.\end{dcases*}$ (14)
Theorem 1.
$A^{*}$ is the unique solution to the $l_{1}$-norm estimator (11) at time $T$ if
$\sum_{t=0}^{T-1}z_{t}^{i}(y)>0,\quad\forall y\in\mathbb{S}^{d-1}$ (15)
holds for all $i\in\{1,\dots,d\}$.
Proof.
The equivalent condition for $A^{*}$ to be the unique solution of the convex optimization problem (13) is the existence of some $\epsilon>0$ such that
$\sum_{t=0}^{T-1}\|w_{t}\|_{1}<\sum_{t=0}^{T-1}\|\Delta\cdot x_{t}+w_{t}\|_{1},$
$\hskip 96.73936pt\forall\Delta\in\mathbb{R}^{d\times d}:0<\|\Delta\|_{F}\leq\epsilon,$ (16)
since a strict local minimum in convex problems implies the unique global minimum. A sufficient condition for (IV-A) is to satisfy all coordinate-wise inequalities. That is, if there exist $\epsilon_{1},\dots,\epsilon_{d}>0$ such that
$\sum_{t=0}^{T-1}|w_{t}^{i}|<\sum_{t=0}^{T-1}|\Delta_{i}^{T}x_{t}+w_{t}^{i}|,\quad\forall\Delta_{i}\in\mathbb{R}^{d}:0<\|\Delta_{i}\|_{2}\leq\epsilon_{i}$ (17)
for all $i\in\{1,\dots,d\}$, then the inequality (IV-A) is satisfied. For all $i$, consider a sufficiently small $\epsilon_{i}>0$. Then, we have
$|\Delta_{i}^{T}x_{t}+w_{t}^{i}|=(\Delta_{i}^{T}x_{t}+w_{t}^{i})\cdot\text{sgn}(w_{t}^{i})=\Delta_{i}^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})+|w_{t}^{i}|$
for $w_{t}^{i}\neq 0$. Substituting the above equation into the right-hand side of (17) yields
$\sum_{t=0}^{T-1}|w_{t}^{i}|<\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}(\Delta_{i}^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})+|w_{t}^{i}|)+\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}=0\end{subarray}}^{T-1}|\Delta_{i}^{T}x_{t}|,$
which is simplified to
$0<\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}(\Delta_{i}^{T}x_{t}\cdot\text{sgn}(w_{t}^{i}))+\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}=0\end{subarray}}^{T-1}|\Delta_{i}^{T}x_{t}|.$
for all $0<\|\Delta_{i}\|_{2}\leq\epsilon_{i}$. For all $i$, dividing both sides by $\|\Delta_{i}\|_{2}>0$ leads to the set of inequalities in (15).
To attain the exact recovery, it suffices to show that the random variables on the left-hand sides of (15) are sufficiently positive with high probability. In the following two subsections, each will present a distinct scenario in which the $l_{1}$-norm estimator achieves the exact recovery despite the presence of adversarial noncentral disturbances.
@@ IV-B Exact Recovery under Sign-restricted Disturbances
In this subsection, we will demonstrate that the exact recovery is achieved under the condition that each disturbance has equal chance of being positive or negative. The structure of the disturbance $w_{t}$ consists of $\alpha_{t}$ and $\beta_{t}$, where $\alpha_{t}$ is a sub-Gaussian vector symmetric around zero and $\beta_{t}>0$ is a scaling factor. The formal assumption is given below.
Assumption 5 (Sign-restricted disturbances):
Given the filtration $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$, a disturbance $w_{t}\in\mathbb{R}^{d}$ at the attack times is formulated as
@@
  • $\beta_{t}$ is a scaling random vector conditioned on $\mathcal{F}_{t}$, where $0<\beta_{t}^{i}<\infty$ holds for all $i=1,\dots,d$.
@@
Remark 1 (Necessity of symmetry assumption):
To formulate adversarial noncentral disturbances, we have introduced $\alpha_{t}$ that is symmetric around zero. We adopted this assumption since the system cannot defend the attack if the adversary has a full information set $\mathcal{F}_{t}$ and the freedom of choosing any nonzero-mean attack. For example, consider the scalar system $x_{t+1}=a^{*}x_{t}+w_{t}$, where the true system satisfies $10<a^{*}<$ 1 and $10<x_{0}<$ 1. Suppose that $w_{t}$ conditioned on $\mathcal{F}_{t}$ is chosen by an adversary as $-\text{sgn}(x_{t})$ for all $t\geq 0$. In such a case, one always attains
@@
However, any valid estimation method tries to minimize the distance between $x_{t+1}$ and $\hat{a}x_{t}$ to find an estimate $\hat{a}$. One can thus only arrive at a negative estimate $\hat{a}$ despite the true system $a^{*}>0$, since we have $\text{sgn}(x_{t})\neq\text{sgn}(x_{t+1})$ for all $t\geq 0$. This occurs since the attacker can always deceive the system if the attack can be universally unidirectional based on the filtration. The symmetry assumption (19) on $\alpha_{t}$ prevents such a phenomenon and enables system identification.
The disturbance structure (18) includes zero-mean Gaussian variables since $\alpha_{t}$ can be zero-mean Gaussian and $\beta_{t}$ can be a constant random vector. Our formulation will later generalize this by converting a symmetric distribution into a noncentral distribution by adjusting a scaling factor $\beta_{t}$.
Now, we analyze the sufficient conditions for the exact recovery shown in Theorem 1. We first provide the following lemma given in [zhang2024exact] for the upper term in (14).
Lemma 3.
Let $N_{T}$ be the cardinality of the set $\{1\leq t\leq T-1:w_{t}\equiv 0,w_{t-1}\in\mathcal{K}_{T}\}$, which is the number of non-attack times immediately preceded by an attack. Suppose that Assumptions 2, 3, and 4 hold. Given $y\in\mathbb{S}^{d-1}$ and $i\in\{1,\dots,d\}$, we have
$\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}=0\end{subarray}}^{T-1}|y^{T}x_{t}|\geq\frac{c\lambda^{5}N_{T}}{\sigma_{w}^{4}}$ (20)
with probability $1-\exp\bigr{(}-\Theta\bigr{(}\frac{\lambda^{4}N_{T}}{\sigma_{w}^{4}}\bigr{)}\bigr{)}$, where $c$ is a positive absolute constant.
Now, given $y\in\mathbb{S}^{d-1}$, we present the analysis on bounding the tail probability of the left-hand side of (15), for which we will study the sub-Gaussian norms under Assumption 5.
Lemma 4.
Suppose that Assumptions 1, 2, 3, 4, and 5 hold. Define $\tau:=\sigma_{w}/\lambda$. Given $y\in\mathbb{S}^{d-1}$, $i\in\{1,\dots,d\}$, and $\delta\in(0,1]$, the condition (15) holds when
$T\geq\Theta\biggr{(}\frac{\max\{1,\tau^{10}\}}{p(1-p)^{2}(1-\|A^{*}\|_{2})^{2}}\cdot\log\biggr{(}\frac{1}{\delta}\biggr{)}\biggr{)}$ (21)
with probability at least $1-\delta$.
Proof.
We first analyze the sub-Gaussian parameter of the lower term in (14). Note that the symmetry assumption (19) implies that the adversarial disturbance $w_{t}$ also satisfies
@@
for all $\mathcal{F}_{t}$ due to $\beta_{t}>0$. In turn, at the attack times, we have $\mathbb{E}[y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})]=\mathbb{E}[y^{T}x_{t}\cdot\mathbb{E}[\text{sgn}(w_{t}^{i})~{}|~{}\mathcal{F}_{t}]]=0$. Thus, we can leverage Definition 2 to derive the sub-Gaussian parameter of the relevant term. Under Assumption 3, the state equation is written as
$x_{t}=(A^{*})^{t}x_{0}+\sum_{k\in[0,t-1]\cap\mathcal{K}_{T}}(A^{*})^{t-1-k}w_{k}$
due to the system dynamics (8). Define another filtration
$\mathcal{F}^{i}=\bm{\sigma}\{\text{sgn}(w_{t}^{i}):0\leq t\leq T-1,w_{t}^{i}\neq 0\}.$
Then, for all $\lambda\in\mathbb{R}$ and for all $s\in[0,T-2]\cap\mathcal{K}_{T}$, we have
@@
$\mathbb{E}\Bigr{[}\exp\bigr{(}\lambda\sum_{\begin{subarray}{c}t=s+1,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}(A^{*})^{t-1-s}w_{s}\cdot\text{sgn}(w_{t}^{i})\bigr{)}~{}\Bigr{|}~{}\mathcal{F}_{s},\mathcal{F}^{i}\Bigr{]}$
@@
$\hskip 82.51299pt\leq\exp\biggr{(}\lambda^{2}\cdot\Theta\Bigr{(}\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$ (22)
due to
$\Biggr{\|}\sum_{\begin{subarray}{c}t=s+1,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}(A^{*})^{t-1-s}w_{s}\cdot\text{sgn}(w_{t}^{i})\Biggr{\|}_{\psi_{2}}$
$\hskip 28.45274pt\leq\sum_{\begin{subarray}{c}t=s+1,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}\bigr{\|}(A^{*})^{t-1-s}w_{s}\cdot\text{sgn}(w_{t}^{i})\bigr{\|}_{\psi_{2}}$
$\hskip 28.45274pt\leq\sum_{\begin{subarray}{c}t=s+1,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}\|A^{*}\|_{2}^{t-1-s}\|w_{s}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}},$ (23)
conditioned on $\mathcal{F}_{s}$ and $\mathcal{F}^{i}$. The first inequality comes from Lemma 2, the triangle inequality, and $\|y\|_{2}=1$. The second inequality is due to Lemma 2 and the fact that alternating the sign does not affect the sub-Gaussian norm since the definition (2) involves the squared variable. From (23), the equivalence of Definitions 1 and 2 yields (22).
By repeatedly selecting from the largest to the smallest element for $s\in[0,T-2]\cap\mathcal{K}_{T}$, applying the tower rule conditioning on $(\mathcal{F}_{s},\mathcal{F}^{i})$, and leveraging (22), we obtain
$\mathbb{E}\Biggr{[}\exp\biggr{(}\lambda\biggr{[}\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}x_{t}\cdot\text{sgn}(w_{t}^{i})\biggr{]}\biggr{)}\Biggr{]}$
@@
$=\mathbb{E}\Biggr{[}\exp\biggr{(}\lambda\biggr{[}\sum_{\begin{subarray}{c}t=0,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}\Bigr{[}y^{T}(A^{*})^{t}x_{0}\cdot\text{sgn}(w_{t}^{i})\Bigr{]}$
$\hskip 17.07164pt+\sum_{k\in[0,T-2]\cap\mathcal{K}_{T}}\sum_{\begin{subarray}{c}t=k+1,\\ w_{t}^{i}\neq 0\end{subarray}}^{T-1}y^{T}(A^{*})^{t-1-k}w_{k}\cdot\text{sgn}(w_{t}^{i})\Bigr{]}\biggr{]}\biggr{)}\Biggr{]}$